OpenStudy (yolo_king):

If A is located at (-3,5) and B is located at (3,-2), find point C that is 4/5 the distance from point A to point B.

OpenStudy (yolo_king):

How do I solve this?

OpenStudy (aum):

Find the x-distance between the two points and find a point 4/5th in the x-direction. Find the y-distance between the two points and find a point 4/5th in the y-direction. Combine the above two and put it in coordinate form: (x, y).

OpenStudy (aum):

A is (-3, 5) B is (3, -2) x-distance between A and B is: 3 - (-3) = 6. y-distance between A and B is: -2 - (5) = -7. 4/5 x 6 = 24/5 = 4.8 4/5 x (-7) = -28/5 = -5.6 4.8 is the x-distance between points A and C. -5.6 is the y-distance between points A and C. What is the x-coordinate of C? What is the y-coordinate of C?

OpenStudy (aum):

Add the x-distance to the x-coordinate of A. Add the y-distance to the y-coordinate of A. That will give you the x and y coordinates of C.

OpenStudy (yolo_king):

Ahhhhh ok thank you so much for the explanation. I get frustrated when my teachers rush through explaining these complex assignments and C = -0.8 THanks A lot :))

OpenStudy (aum):

You are welcome. x-coordinate of C is: x-coordinate of A + x-distance between A and C = -3 + 4.8 = 1.8. y-coordinate of C is: y-coordinate of A + y-distance between A and C = 5 - 5.6 = -0.6. C is at (1.8, -0.6)